# Debiased Machine Learning and Network Cohesion for Doubly-Robust Differential Reward Models in Contextual Bandits

### v2

@@ 4.2 Regret Analysis
Under Assumptions 1 and 2, with probability at least $1-\delta$, Algorithm 1’s regret satisfies
$\left(\beta_{K}(\delta^{\prime})+\gamma_{K}(\delta^{\prime})\left[1+\frac{4}{d}\right]\right)\sqrt{4cH_{K}Kd\log\left(\gamma+\lambda M+\frac{K+1}{8d}\right)-\log\det\left(V_{0}\right)}$
$+\frac{4\gamma_{K}(\delta^{\prime})}{d}\sqrt{\frac{8K}{\lambda}\log\left(\frac{4}{\delta}\right)},$
where $f_{K}=\operatorname{logdet}V_{K}-\operatorname{logdet}V_{0}$, $V_{0}=\lambda L\otimes I_{p}+\gamma I_{np}$, $\delta^{\prime}=\delta/4K$, $\min(\pi(0|s),1-\pi(0|s))>1/c$ and $H_{K}=O(log(K))$ is the harmonic number. $\beta_{K}$ and $\gamma_{K}$ are defined in Appendix G.
Proof of Theorem 1 is in Appendix G. The regret bound is similar to prior work by Abeille & Lazaric (2017); however, our bound differs in three ways: (1) the harmonic number $H_{K}$ enters as an additional cost for considering average regret per stage with the regret being $O(\sqrt{K}\log^{2}(K))$ having an additional $\log(K)$ factor; (2) the bound depends on the dimension of the differential reward model rather than the dimension of the overall model which can significantly improve the regret bound; and (3) the main benefit of our use of DML is in $\beta_{K}(\delta^{\prime})$ and $\gamma_{K}(\delta^{\prime})$, which depend on the rate of convergence of the model $f$ to the true mean differential reward $r$ as discussed in Appendix G.2, which demonstrates the benefits of good models for this term and how it impacts regret.
Note this regret bound is sublinear in the number of stages and scales only with the differential reward complexity $d$, not the complexity of the baseline reward $g$. As the second term scales with $\sqrt{K}$, for large $K$ we can focus on the first term which scales $O\left(\sqrt{c\cdot d\cdot\log^{2}(K)K}\right)$. Prior work scales sublinearly with the number of decision times $T$ as they assume either a single contextual bandit ($n=1$) or a fixed number of individuals $n$. Greenewald et al. (2017) scales as $O(d^{2}\sqrt{T\log(T)})$ while Yang et al. (2020) scale as $O(\sqrt{\tilde{d}nT\log(T)})$ where $\tilde{d}$ is the complexity of the joint baseline and differential reward model. Interestingly, at stage $K$ we see $n=K$ individuals over $n=K$ decision times (with different number of observations per individual); however, we do not see a regret scale with $\sqrt{nT}=K$. Instead, we only receive an extra $\log(K)$ factor reflecting the benefit of pooling on average regret.
There are several technical challenges to this regret bound. First, the confidence ellipsoids $\beta_{k}$ and $\gamma_{k}$ depend on the sub-Gaussian variance factor of the pseudo-reward and need to be derived for the DML pseudo-reward. Second, two results in Abbasi-Yadkori et al. (2011) need to be reproven: the first is Lemma 7, a linear predictor bound that is used as a key step to derive the final regret bound. It requires handling the fact that our regularized least squares estimate now uses our DML pseudo-reward instead of the observed reward. The second, Proposition 1, requires care as we are now doing weighted regularized least squares. The original version used in Abeille & Lazaric (2017) requires an upper bound on the sum of squared feature norms, but applying Abbasi-Yadkori et al. (2011) only gives us an upper bound on the sum of weighted squared norms. In order to derive the needed bound, we use assumption 4 and then apply Abbasi-Yadkori et al. (2011) to obtain our version of their results that has an upper bound that depends on $c>0$, the lower bound on the weights. Finally, the regret bound itself needs to handle the fact that we have stages with multiple individuals (increasing by one) per stage. This leads to a sum over stages and participants within stages of the difference between RLS and TS (and RLS and true) linear predictors. By some manipulation and an application of Cauchy Schwartz, we see a sum of $\frac{1}{k}$ over stages, which leads to the harmonic number, which describes the additional cost of handling multiple participants in a study in stages.
@@ 5.1 Competitor Comparison Simulation
In this section, we test three versions of our proposed method: (1) DML-TS-NNR-BLM: Our algorithm using an ensemble of Bagged Linear Models, (2) DML-TS-NNR-BT: Our algorithm using an ensemble of Bagged stochastic gradient Trees (Gouk et al., 2019; Mastelini et al., 2021), and (3) DML-TS-SU-BT: Same as (2) but treating the data as if it were derived from a Single User.
@@
We implemented these using Option 2 in Algorithm 1, and compared to four related methods: (1) Standard: Standard Thompson sampling for linear contextual bandits, (2) AC: Action-Centered contextual bandit algorithm (Greenewald et al., 2017), (3) IntelPooling: The intelligentpooling method of Tomkins et al. (2021) fixing the variance parameters close to their true values, and (4) Neural-Linear: a method that uses a pre-trained neural network to transform the feature space for the baseline reward (similar to the Neural Linear method of Riquelme et al. (2018), which in turn was inspired by Snoek et al. (2015)). In general, we expect our method to outperform these methods because it is the only one that can (1) efficiently pool across users and time, (2) leverage network information, and (3) accurately model a complex, nonlinear baseline reward.
@@
We compare these seven methods under three settings that we label as Homogeneous Users, Heterogeneous Users, and Nonlinear. The first two settings involve a linear baseline model and time-homogeneous parameters, but they differ in that the users in the second setting have distinct parameters. The third setting is more general and includes a nonlinear baseline, user-specific parameters, and time-specific parameters. Across all three settings, we simulate 125 stages following the staged recruitment regime depicted in Figure 3 in Appendix A, and we repeat the full 125-stage simulation 50 times. Appendix D provides details on the setup and a link to our implementation.
Figure 1: Cumulative regret in the (a) Homogeneous Users, (b) Heterogeneous Users, and (c) Nonlinear settings. The DML methods perform competitively in all three settings and appear to be achieving sublinear regret as expected based on our theoretical results. The DML-TS-NNR-BLM and DML-TS-NNR-BT algorithms perform best, and their final regret is statistically indistinguishable (see Table 1 in Appendix D.2).
Figure 1 shows the cumulative regret for each method at varying stages. DML methods perform competitively against the benchmark methods in all three settings and achieve sublinear regret as expected based on our theoretical results. Across all settings, the best-performing method is either DML-TS-NNR-BLM or DML-TS-NNR-BT. In the first setting, the difference between our methods and IntelPooling is not statistically meaningful because IntelPooling is properly specified and network information is not relevant. In the other two settings, our methods offer substantial and statistically meaningful improvement over the other methods. Appendices D.2 and D.3 shows detailed pairwise comparisons between methods and an additional simulation study using a rectangular array of data.
5.2 Valentine Results
In parallel with the simulation study, we conducted a comparative analysis on a subset of participants from the Valentine Study (Jeganathan et al., 2022), a prospective, randomized-controlled, remotely-administered trial designed to evaluate an mHealth intervention to supplement cardiac rehabilitation for low- and moderate-risk patients. In the analyzed subset, participants were randomized to receive or not receive contextually tailored notifications promoting low-level physical activity and exercise throughout the day. The six algorithms being compared include (1) Standard, (2) AC, (3) IntelPooling, (4) Neural-Linear, (5) DML-TS-SU-RF (RF stands for Random Forest (Breiman, 2001)), and (6) DML-TS-NNR-RF. Figure 2 shows the estimated improvement in average reward over the original constant randomization, averaged over stages (K = 120) and participants (N=108).
To demonstrate the advantage of our proposed algorithm in terms of average reward compared to the competing algorithms, we conducted a pairwise paired t-test with a one-sided alternative hypothesis. The null hypothesis ($H_{0}$) stated that two algorithms achieve the same average reward, while the alternative hypothesis ($H_{1}$) suggested that the column-indexed algorithm achieves a higher average reward than the row-indexed algorithm. Figure 2 displays the p-values obtained from these pairwise t-tests. Since the alternative hypothesis is one-sided, the resulting heatmap is not symmetric. More details on implementation can be found in Appendix E.
@@ 5.2 Valentine Results
Figure 2: (left) Boxplot of unbiased estimates of the average per-trial reward for all six competing algorithms, relative to the reward obtained under the pre-specified Valentine randomization policy across 100 bootstrap samples. Within each box, the asterisk ($\ast$) indicates the mean value, while the mid-bar represents the median. (right) Heatmap of p-values from the pairwise paired t-tests. The last column’s dark shade indicates that the proposed DML-TS-NNR-RF algorithm achieves significantly higher rewards than the other five competing algorithms. This implies that after implementing our proposed algorithm, the step counts increased by $3.5\%$ more than what the constant randomization policy achieved.
@@
To further enhance the competitive performance of our proposed DML-TS-NNR algorithm, we conducted an additional comparative analysis using a real-world dataset from the Intern Health Study (IHS) (NeCamp et al., 2020). Further details regarding the analysis can be found in Appendix F.
6 Discussion and Future Work
In this paper, we have presented the DML Thompson Sampling with Nearest Neighbor Regularization (DML-TS-NNR) algorithm, a novel contextual bandit algorithm specifically tailored to the mHealth setting. By leveraging the DML framework and network cohesion penalties, DML-TS-NNR is able to accurately model complex, nonlinear baseline rewards and efficiently pool across both individuals and time. The end result is increased statistical precision and, consequently, the ability to learn effective, contextually-tailored mHealth intervention policies at an accelerated pace.
While DML-TS-NNR achieves superior performance relative to existing methods, we see several avenues for improvement. First, the algorithm considers only immediate rewards and, as such, may not adequately address the issue of treatment fatigue. Second, the current algorithm involves computing a log-determinant and matrix inverse, which can be computationally expensive for large matrices. Third, we have made the simplifying assumption that the differential reward is linear in the context vectors. Fourth, we have assumed that the network structure is known and contains only binary edges. Fifth, our algorithm involves several hyperparameters whose values may be difficult to specify in advance. Future work will aim to address these practical challenges in applied settings.
@@ Appendix A Recruitment Regime Illustration
Appendix B Additional Details for Algorithm 1
The inclusion of $\gamma>0$ ensures a unique solution at each step, while $\lambda$ controls smoothness across the individual-specific and time-specific parameters. In particular, when $\lambda=0$, no information is shared across individuals or across time. A positive value of $\lambda$ introduces the smoothness of these estimates, e.g., if $(i,j)\subset E_{\text{ind}}$ then $\|\theta_{i}^{\text{ind}}-\theta_{j}^{\text{ind}}\|_{2}$ tends to be small. It can be shown that $\sum_{(i,j)\subset E_{\text{ind}}}\|\theta_{i}^{\text{ind}}-\theta_{j}^{\text{ind}}\|_{2}^{2}=\theta_{\text{ind}}^{\top}L_{\otimes}\theta_{\text{ind}}$, where $L_{\otimes}=L\otimes I_{p}$ and $\theta_{\text{ind}}=(\theta_{1}^{\text{ind}},\ldots,\theta_{k}^{\text{ind}})$ is the vector of individual-specific parameters. Let $\boldsymbol{x}_{i,t}:=\boldsymbol{x}(S_{i,t},A_{i,t})$ and $V_{0}=\operatorname{diag}(\gamma I_{p},\lambda L^{\text{ind}}_{\otimes}+\gamma I_{kp},\lambda L^{\text{time}}_{\otimes}+\gamma I_{kp})$. Let $\Phi_{k}$ be the design matrix of features $\phi(\textbf{x}_{i,t})$, $W_{k}$ be the diagonal matrix of weights $\tilde{\sigma}_{i,t}^{2}$, and $R_{k}$ the vector of pseudo-rewards $\tilde{R}_{i,t}^{f^{(k)}}$. In all of these $(i,t)\in\mathcal{O}_{k}$. Then the minimizer $\hat{\Theta}_{k}$ of (5) is
$[\Phi_{k}^{\top}W_{k}\Phi_{k}+V_{0}]^{-1}[\Phi^{\top}_{k}W_{k}R_{k}]$ $=\underbrace{\Big{[}\sum_{(i,t)\in\mathcal{O}_{k}}\tilde{\sigma}_{i,t}^{2}\phi(\boldsymbol{x}_{i,t})\phi(\boldsymbol{x}_{i,t})^{\top}+V_{0}\Big{]}^{-1}}_{V_{k-1}^{-1}}\underbrace{\Big{[}\sum_{(i,t)\in\mathcal{O}_{k}}\tilde{\sigma}^{2}_{i,t}\tilde{R}_{i,t}^{f^{(k)}}\phi(\boldsymbol{x}_{i,t})\Big{]}}_{b_{k-1}}.$
where $\phi(\boldsymbol{x}_{i,t})\in\mathbb{R}^{p(2k+1)}$ for any $(i,t)\in\mathcal{O}_{k}$ is defined as
$\begin{split}\phi(\boldsymbol{x}_{i,t})=(\boldsymbol{x}_{i,t},&\bar{\bf 0}_{(i-1)p},\boldsymbol{x}_{i,t},\bar{\bf 0}_{(k-i)p},\bar{\bf 0}_{(t-1)p},\boldsymbol{x}_{i,t},\bar{\bf 0}_{(k-t)p}).\end{split}$
The first location of non-zero entries is for the global parameters, the second for the individual parameters and the third for the time parameters.
Remark 1 (Computationally efficient estimation).
Direct calculation of $\hat{\Theta}_{k}$ leads to a computationally expensive inversion of the $(2k+1)\cdot p$ dimensional matrix $V_{t-1}$ at each stage $k$. To avoid this, we observe $V_{k}=V_{k-1}+\sum_{(i,t)\in\mathcal{O}_{k}\backslash\mathcal{O}_{k-1}}\tilde{\sigma}_{i,t}^{2}\phi(\boldsymbol{x}_{i,t})\phi(\boldsymbol{x}_{i,t})^{\top}$ and apply the Sherman-Morrison formula for more efficient computation.
Next, we consider a generic randomized algorithm based on the RLS estimate by sampling a perturbed parameter $\tilde{\Theta}_{k}$ and then selecting an action by simply maximizing the linear differential reward $\boldsymbol{x}(s,a)^{\top}(\tilde{\theta}+\tilde{\theta}_{i}+\tilde{\theta}_{t})$. This construction includes standard Thompson sampling as an important special case. Specifically, we construct $\tilde{\Theta}_{k}=\hat{\Theta}_{k}+\beta_{k}(\delta^{\prime})V_{k}^{-1/2}\eta_{k}$ where $\eta_{k}$ is a random sample drawn i.i.d. from a suitable multivariate distribution $\mathcal{D}^{\text{TS}}$ and $\beta_{k}(\delta^{\prime})$ is a term from the self-normalization bound (Theorem 2) developed in Abbasi-Yadkori et al. (2011). Because of our use of $f_{i,t}$ to approximate $r_{i,t}$, $\beta_{k}(\delta^{\prime})$ is smaller, and thus our distribution for $\hat{\Theta}_{k}$ has lower variance than if we did not use $f_{i,t}$. We discuss this in detail in Appendix G.2. To ensure small regret, we must choose the random variable $\eta_{k}$ to ensure sufficient exploration but not too much. Definition 1 is from Abeille & Lazaric (2017) and formalizes the properties of the random variable $\eta_{k}$.
Definition 1.
While a Gaussian prior satisfies Definition 1, this approach allows us to move beyond Bayesian posteriors to generic randomized policies. In practice, the true parameter values $\Theta$ are unknown, so $\beta_{k}(\delta^{\prime})$ is not available. Thus one needs to insert upper bounds for $\|\mathcal{L}^{\text{ind}}_{\otimes}\theta_{\text{ind}}\|_{2}$ and $\|\theta_{\text{ind}}\|_{2}$ and similarly for time-specific parameters. For example, $\|L_{\otimes}\theta\|_{2}\leq M\cdot N\max_{(i,j)\in E}\|\theta_{i}-\theta_{j}\|_{2}$. Similarly, based on $\|\theta\|_{2}\leq\sqrt{k}\max_{i}\|\theta_{i}\|_{2}$, by Assumption 2, we have $\|\theta\|_{2}\leq\sqrt{k}B$.
Appendix C Option 1 & 2
In Algorithm 1, we provide two options for constructing $\hat{f}$ and the corresponding pseudo-outcomes $\tilde{R}_{i,t}^{\hat{f}}$. Option 1 assumes i.i.d. additive error and recomputes the pseudo-rewards, $\tilde{R}_{i,t}^{\hat{f}^{(m)}}$, at each stage for all $i,t$ using an updated estimate of $f$, which can be estimated using either (a) sample splitting or (b) stable estimators, such as bagged ensembles constructed via sub-sampling. Option 2 uses historical data to generate predictions in an online fashion, and calculates the pseudo-rewards only once without updating them for all subsequent stages.
Our simulations have confirmed that both options result in comparable regret. In reality, however, it’s important to note that the decisions made for individual $i$ at time $t$ may depend on all previous data. Based on this argument, Option 2 was used in both case studies and simulations, highlighting the benefits of our approach in numerical as well as real-world mHealth studies, as detailed in the manuscript.
C.1 Option 1
Sample splitting has been widely used in the DML-related literature (Chernozhukov et al., 2018; Kennedy, 2020) to relax modeling constraints in constructing $\hat{f}$. This approach allows the flexibility of the model to increase with the sample size while protecting against overfitting. As a result, complex machine learning algorithms can be utilized to estimate the function $f$, resulting in accurate estimation of the differential reward $\Delta^{f}$ and, consequently, precise estimates of the $\theta$’s in the linear differential reward function.
When implemented as part of our proposed algorithm, sample splitting randomly partitions all available data into $M$ folds. This step relies heavily on the assumption of i.i.d additive errors, as the random splits are formed on a per-observation basis. Theoretical findings outlined in Chernozhukov et al. (2018) demonstrate the crucial role of sample splitting in achieving $\sqrt{n}$-consistency and ensuring the validity of inferential statements for the parameters in the linear differential reward function.
As an alternative to sample splitting, one can instead construct $\hat{f}$ via stable estimators that exhibit a $o(n^{-1/2})$ leave-one-out stability. A recent study by Chen et al. (2022) demonstrated that estimators based on predictive models that satisfy this condition can achieve $\sqrt{n}$-consistency and asymptotic normality without relying on the Donsker property or employing sample splitting.
The stability conditions outlined in Theorem 5 of Chen et al. (2022) are satisfied by bagging estimators formed with sub-sampling. We leverage this result in Section 5 by testing several versions of our method based on bagged estimators: two based on bagged stochastic gradient trees (DML-TS-NNR-BT and DML-TS-SU-BT) and one based on bagged linear models (DML-TS-NNR-BLM).
As long as the errors are i.i.d., methods based on either (a) sample splitting or (b) stable estimators are permissible within our method. In both cases, we are able to leverage both current and previous observations to construct the latest estimation of $\hat{f}$.
C.2 Option 2
In the mobile health setting, deviations from the expected reward typically represent the effects of idiosyncratic error as opposed to adversarial actions. Consequently, the i.i.d. assumption is more plausible in mobile health than in general contextual bandit settings. However, due to the sequential design of mobile health studies, one challenge that may arise is dependence across time within individual users. Option 2 adapts our method to address this challenge.
When considering errors that exhibit temporal dependence, it is reasonable to utilize past observations to create future pseudo-rewards but not the other way around. For instance, at stage $k$, we train the model $\hat{f}^{(k)}$ using all the history observed up until stage $k-1$ and construct pseudo-rewards for the rewards observed in stage $k$; however, we do not use $\hat{f}^{(k)}$ to update pseudo-rewards for previous stages.
Fixing the pseudo-rewards in this manner enables us to move beyond the assumption of i.i.d. additive errors. This approach is compatible with both the sample splitting and bagging approaches discussed previously. As a computational convenience, researchers may choose to update $\hat{f}$ in an online fashion as we did in Section 5.1. The primary tradeoff in doing so is that the online predictive model may not perform as well as a model trained in one batch, especially in the early stages.
Appendix D Additional Details for Simulation Study
The code for the simulation study is fully containerized and publicly available at https://redacted/for/anonymous/peer/review.
D.1 Setup Details
We consider a generative model of the following form for user $i$ at time $t$:
$R_{it}=g(S_{it})+x(S_{it},A_{it})\,\theta_{it}+\epsilon_{it},\quad\epsilon_{it}\sim\mathcal{N}(0,1)$

### v3

@@ 5.2 Regret bound
Under Assumptions 16, with probability at least $1-\delta$, $\boldsymbol{\operatorname{Regret}}_{K}$ satisfies
$O\left(\left(\beta_{K}(\delta^{\prime})+\gamma_{K}(\delta^{\prime})\left[1+\frac{4}{p}\right]\right)\boldsymbol{\cdot}\log(K)\sqrt{4cKp}\right)$
where $\delta^{\prime}=\delta/4K$ and $\min(\pi(0|s),1-\pi(0|s))>1/c$. The exact regret bound is given and both $\beta_{K}$ and $\gamma_{K}$ are defined in Appendix F.
Theorem 1’s proof is in Appendix F. It resembles  Abeille & Lazaric (2017), albeit with adjustments to address (a) the nonlinear baseline and (b) the increasing pool of users and time points. The non-linear baseline requires deriving the new sub-Gaussian variance factor for the centered pseudo-reward and adapting classic contextual bandit theory (Abbasi-Yadkori et al., 2011) to our weighted least squares setting. The increasing pool of users and time points requires modifying the Abeille & Lazaric (2017) proof with a careful use of stages and Cauchy Schwarz. The overall bound scales as
$O\left(p^{3/2}\sqrt{\log^{2}(K)K}\right).$
DML is particularly advantageous for $\beta_{K}(\delta^{\prime})$ and $\gamma_{K}(\delta^{\prime})$, which rely on the convergence rate of the model $f$ towards the true mean differential reward $r$, as discussed further in Appendix F.2. Another interesting aspect of the regret bound is its dependence solely on the dimension, $p$, of the differential reward model, rather than the dimension of the baseline reward. This characteristic allows us to achieve robust regret bounds even in cases where the baseline reward is highly complex.
@@ 6 Experiments
This section presents results from applying DML-TS-NNR in a simulated mHealth study with staggered recruitment and an off-policy comparison study for cardiac rehabilitation mHealth interventions. The simulations were implemented using Python and the results were generated using individual compute nodes with two 3.0 GHz Intel Xeon Gold 6154 processors and 180 GB of RAM. Case studies were implemented using R 4.2.2 and results were generated on a cluster composed of individual compute nodes with 2.10 GHz Intel Xeon Gold 6230 processors and 192 GB of RAM.
@@ 6.1 Competitor Comparison Simulation
In this section, we compare DML-TS-NNR to four competing methods in simulation. We implemented DML-TS-NNR using Option 2 with a bagged ensemble of stochastic gradient trees (Gouk et al., 2019; Mastelini et al., 2021) trained online via the River library (Montiel et al., 2021). We programmed the linear algebra operations using SuiteSparse (Davis & Hu, 2011) to take advantage of the sparse nature of the Laplacian matrices and $\phi_{i,t}$ from Algorithm 2.
@@
Our competing baselines are (a) Standard: Standard Thompson sampling for linear contextual bandits, (b) AC: the Action-Centered contextual bandit algorithm (Greenewald et al., 2017), (c) IntelPooling: The intelligentpooling method of Tomkins et al. (2021) fixing the variance parameters close to their true values, and (d) Neural-Linear: a method that uses a pre-trained neural network to transform the feature space for the baseline reward (similar to the Neural Linear method of Riquelme et al. (2018)).
@@
We compare these methods under three settings: Homogeneous Users, Heterogeneous Users, and Nonlinear. The first two involve a linear baseline model and time-homogeneous parameters, with the second setting having distinct user parameters. The third setting, designed to mirror the challenges of mHealth studies, includes a nonlinear baseline and both user- and time-specific parameters. For each setting, we simulate 200 stages following the staged recruitment regime depicted in Figure 3 in Appendix A, and we repeat the full 200-stage simulation 50 times. Appendix B.1 provides details on the setup and a link to our implementation.
Figure 1: Cumulative regret in the (a) Homogeneous Users, (b) Heterogeneous Users, and (c) Nonlinear settings. DML-TS-NNR performs competitively in the first setting (the simplest), and it substantially outperforms the next-best method (IntelPooling) in the others. Pairwise comparisons in Table 2 in Appendix B.3) reveal that DML-TS-NNR outperformed all competitors shown above in 49 of 50 repetitions in the Nonlinear setting (panel (c)).
Figure 1 displays the average cumulative regret for each method. DML-TS-NNR and IntelPooling have similar performance in the Homogeneous Users setting, but DML-TS-NNR excels in the Heterogeneous Users and Nonlinear settings, with significantly lower cumulative regret by stage 200. This is expected as DML-TS-NNR can utilize network information and model nonlinear baselines in these settings. However, in the Homogeneous Users setting, where no network information is available and a linear baseline is used, DML-TS-NNR does not outperform IntelPooling.
Appendices B.2, B.3, and B.4 offer more results, including other method comparisons, a rectangular data array simulation, and pairwise statistical comparisons. The latter show DML-TS-NNR outperformed all competitors in at least 49 of 50 repetitions in the Nonlinear setting, indicating even higher statistical confidence than Figure 1 suggests.
6.2 Valentine Study Analysis Results
@@
Figure 2: (left) Boxplot of unbiased estimates of the average per-trial reward for all five competing algorithms, relative to the reward obtained under the pre-specified Valentine randomization policy across 100 bootstrap samples. Within each box, the asterisk ($\ast$) indicates the mean value, while the mid-bar represents the median. (right) Heatmap of p-values from the pairwise paired t-tests.
In this section, we compare DML-TS-NNR to the above algorithms via off-policy evaluation using data from the Valentine Study (Jeganathan et al., 2022), a prospective, randomized-controlled, remotely-administered trial designed to evaluate an mHealth intervention to supplement cardiac rehabilitation for low- and moderate-risk patients. In the analyzed dataset, participants were randomized to receive or not receive contextually tailored notifications promoting low-level physical activity and exercise throughout the day. The left of Figure 2 shows the estimated improvement in the average reward over the original constant randomization, averaged over stages (K = 120) and participants (N=108). We see that DML-TS-NNR achieved the highest average reward; its average performance (across bootstrap replications) is higher than the 75th percentile of all other methods.
To further analyze the improvement of DML-TS-NNR relative to the other methods, we aim to test whether the proposed algorithm consistently and significantly improves cumulative rewards across all assessed datasets. Here we conducted pairwise paired t-tests with one-sided alternative hypotheses. The null hypothesis ($H_{0}$) for these tests is that the two algorithms being compared achieve the same average reward. The alternative hypothesis ($H_{1}$) is that the algorithm listed in the column achieves higher average rewards than the algorithm listed in the row. The right of Figure 2 displays the p-values obtained from these pairwise t-tests. The last column’s dark shade indicates that the proposed DML-TS-NNR algorithm achieves significantly higher rewards than the other five competing algorithms ($p\leq 0.01$ for all pairwise comparisons).. This implies that after implementing our proposed algorithm, the step counts increased by $3.5\%$ more than what the constant randomization policy achieved. This translates to an increase of 140 steps per day in step count, given a baseline of 1000 steps per hour. See Appendix C for additional implementation details.
@@
We performed an additional off-policy comparison using data from the Intern Health Study (IHS) (NeCamp et al., 2020), as shown in Figure 10. The results further demonstrate the competitive performance of our DML-TS-NNR algorithm, with the AC algorithm also showing comparable performance in this dataset. Further details on the analysis can be found in Appendix D.
7 Discussion
This paper introduces DML-TS-NNR, a new contextual bandit algorithm for mHealth studies, using DML and network cohesion penalties to model complex rewards and pool information across individuals and time. This enhances statistical precision, enabling rapid learning of accurate, context-specific mHealth treatment policies and improved rewards.
While DML-TS-NNR surpasses current methods, it has four potential improvement areas: it only considers immediate rewards, ignoring treatment fatigue; its computation of log-determinants and matrix inverses is expensive for large matrices; it assumes linear differential rewards and a known network with binary edges; and it requires specification of several hyperparameters. Future work will address these issues.
@@ Appendix A Recruitment Regime Illustration
Appendix B Additional Details for Simulation Study
This appendix provides details regarding the setup of our simulation study, additional results not presented in the main paper, and results from additional simulations employing different competitors and/or simulation designs. We summarize the differences between the competing algorithms in Table 1.
User-specific Parameters Time-specific parameters Pools Across Users Enforces User Network Cohesion Enforces Time Network Cohesion Allows Nonlinear Baseline Nonlinear Baseline Model
DML-TS-NNR $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
DML-TS-NNR-BLM $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\times$
DML-TS-SU $\times$ $\surd$ $\surd$ N/A $\surd$ $\surd$ $\surd$
NNR-Linear $\surd$ $\times$ $\surd$ $\surd$ N/A $\times$ $\times$
IntelPooling $\surd$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$
Neural-Linear $\surd$ $\times$ $\times$ N/A N/A $\surd$ $\surd$
Standard $\surd$ $\times$ $\times$ N/A N/A $\times$ $\times$
AC $\surd$ $\times$ $\times$ N/A N/A $\surd$ $\times$
Table 1: This table compares the methods tested in the simulation study based on the design components listed in the columns, with $\surd$ indicating that the method includes the component listed in the column, $\times$ indicating that it does not, and N/A indicating that the component is not applicable. This table can be used to investigate the effect of certain model components on cumulative regret; for example, comparing DML-TS-NNR and DML-TS-NNR-BLM helps us understand the impact of modeling the nonlinear baseline—the only column that differs between these two methods.
B.1 Setup Details
The code for the simulation study is fully containerized and publicly available at https://github.com/anonymous-conf-submitter-896697/mixed-effects-bandit. The simulation study involves a generative model of the following form for user $i$ at time $t$:
$R_{it}=g(S_{it})+x(S_{it},A_{it})^{{\top}}\,\theta_{it}+\epsilon_{it},\quad\epsilon_{it}\sim\mathcal{N}(0,1)$